# ELOQ: Resources for Enhancing LLM Detection of Out-of-Scope Questions

### v3

@@ 4.3. Semantic Relevance
To demonstrate that the generated out-of-scope questions are indeed semantically similar to the document they are generated from, we conduct retrieval experiments using all out-of-scope questions against our document corpus (2000 document). We employ the following retrieval models: BM25  [^5] https://github.com/castorini/pyserini, SBERT  [^6] https://huggingface.co/sentence-transformers/all-mpnet-base-v2, bge-large-en-v1.5  [^7] https://huggingface.co/BAAI/bge-large-en-v1.5, stella_en_1.5B_v5888https://huggingface.co/NovaSearch/stella_en_1.5B_v5, and Linq-Embed-Mistral [^9] https://huggingface.co/Linq-AI-Research/Linq-Embed-Mistral. Among these models, BM25 and SBERT are well-known retrievers that serve as established baselines, bge-large-en-v1.5 was accepted in SIGIR 2024 resource track as a high-performance embedding model, stella_en_1.5B_v5 ranks #1 on MTEB’s English v1 dataset (Muennighoff et al., 2022), and Linq-Embed-Mistral leads multiple benchmarks, including BEIR (Kamalloo et al., 2023) and MTEB’s English v2 dataset at the time of writing. As shown in Table 6, the Recall@1 values across all models suggest that for approximately half of the out-of-scope questions, the top-retrieved document is the same one from which the question was generated. When we relax this condition to Recall@10, around 80% of out-of-scope questions retrieve their associated document within the top 10 results. This demonstrates the strong semantic connection between out-of-scope questions and their source documents despite the fact that the document lacks the information needed to answer the question.
@@ 4.4. Ablation Study
We examined the impact of various prompts on “defusion detection” task. As seen in Table 7, incorporating examples increased accuracy from 87.61% to 97.35%, and applying self-consistency with $m=3$ further boosted performance to 98.23%. However, further increasing $m$ did not lead to additional improvements. Since ELOQ-Gold is a smaller dataset, using $m=9$ did not yield gains, but this doesn’t rule out its potential benefit on the larger ELOQ-Silver dataset. Therefore, we adopted $m=9$ in our experiments to balance accuracy and expenses (Appendix C).
5. Conclusion and Future Works
We introduced ELOQ, a benchmark for enhancing LLM detection of out-of-scope questions for a given document. Human annotation validates that ELOQ achieves 94.91% accuracy in generating out-of-scope and in-scope questions, while our AutoDefusionEval achieves 98.23% accuracy in detecting effective defusion responses. We further demonstrated ELOQ’s utility by training out-of-scope detectors. For future work, we plan to apply for OpenAI funding to utilize more powerful models, such as o1, to improve both data quality and scale. Additionally, we aim to explore post-training techniques to fine-tune LLMs on ELOQ, ensuring they align better with human preferences by clarifying the confusing aspects of a question rather than generating hallucinated answers.
Appendix A Out-of-scope Example
Document:
Question: How does Justin Jefferson’s exceptional catching ability influence the Vikings’ decision to select him in the first round of the 2020 NFL Draft despite his lack of speed?
Explanation: This question is out-of-scope as it asks about ”Vikings’ decision to select Jefferson in the first round of the 2020 NFL Draft” which was never discussed in the document.
@@ B.1. Extract Claims Prompt
B.2. Recover Missing Claims Prompt
@@ B.3. Out-of-scope Judgement Prompt
B.4. Response Defusion Judgement Prompt
@@ B.8. RAG Zero-shot-CoT
B.9. In-scope Question Generation Prompt
@@
B.10. Out-of-scope Question Generation Prompt
@@ B.11. Remove Claims Prompt
{num_true_fact} ground-truth facts
@@
Appendix C Expenses
Appendix D Embedding Classifier Implementation Details
@@
The unused-token’s hidden state is used as input features for a 2-layer MLP, which is optimized with a BCE loss. A dropout of 0.1 is applied after the first layer. We use Adam optimizer with 1e-4 learning rate. We train each classifier for 10 epochs with batch size 8. The best validation checkpoint is saved and its performance is reported on ELOQ-Gold.
Acknowledgements.

### v4

@@ 4.3. Semantic Relevance
To demonstrate that the generated out-of-scope questions are indeed semantically similar to the document from which they are generated, we conduct retrieval experiments using all out-of-scope questions against our document corpus (2,000 documents). We employ the following retrieval models: BM25  [^5] https://github.com/castorini/pyserini, SBERT  [^6] https://huggingface.co/sentence-transformers/all-mpnet-base-v2, bge-large-en-v1.5  [^7] https://huggingface.co/BAAI/bge-large-en-v1.5, stella_en_1.5B_v5888https://huggingface.co/NovaSearch/stella_en_1.5B_v5, and Linq-Embed-Mistral [^9] https://huggingface.co/Linq-AI-Research/Linq-Embed-Mistral. Among these models, BM25 and SBERT are well-known retrievers that serve as established baselines, bge-large-en-v1.5 was accepted in SIGIR 2024 resource track as a high-performance embedding model, stella_en_1.5B_v5 ranks #1 on MTEB’s English v1 dataset (Muennighoff et al., 2022), and Linq-Embed-Mistral leads multiple benchmarks, including BEIR (Kamalloo et al., 2023) and MTEB’s English v2 dataset at the time of writing. As shown in Table 6, the Recall@1 values across all models suggest that for approximately half of the out-of-scope questions, the top-retrieved document is the same one from which the question was generated. When we relax this condition to Recall@10, around 80% of out-of-scope questions retrieve their associated document within the top 10 results. This demonstrates a strong semantic connection between out-of-scope questions and their source documents, despite the fact that the document lacks the necessary information to answer the question.
@@ 4.4. Ablation Study
We examined the impact of various prompts on “defusion detection” task. As seen in Table 7, incorporating examples increased accuracy from 87.61% to 97.35%, and applying self-consistency with $m=3$ further boosted performance to 98.23%. However, further increasing $m$ did not lead to additional improvements. Since ELOQ-Gold is a smaller dataset, using $m=9$ did not yield gains, but this doesn’t rule out its potential benefit on the larger ELOQ-Silver dataset. Therefore, we adopted $m=9$ in our experiments to balance accuracy and expenses (Appendix A).
5. Conclusion and Future Work
We introduced ELOQ, a benchmark for enhancing LLM detection of out-of-scope questions for a given document. Human annotation validates that ELOQ achieves 94.91% accuracy in generating out-of-scope and in-scope questions, while our AutoDefusionEval achieves 98.23% accuracy in detecting effective defusion responses. We further demonstrated ELOQ’s utility by training out-of-scope detectors, and the results show that a small model (Llama 3.1 8B) can achieve comparable results to the larger one (Llama 3.1 70B) without training. For future work, we plan to apply for OpenAI funding to utilize more powerful models, such as o1, to improve both data quality and scale. Additionally, we aim to explore post-training techniques to fine-tune LLMs on ELOQ, ensuring they align better with human preferences by identifying the confusing aspects of a question rather than generating hallucinated answers.
Acknowledgements.
Appendix
Appendix A Expenses
@@ B.1. Extract Claims
B.2. Recover Missing Claims
@@ B.3. Out-of-scope Judgement
B.4. Response Defusion Judgement
@@ B.8. RAG Zero-shot-CoT
B.9. In-scope Question Generation
@@
B.10. Out-of-scope Question Generation
@@ B.11. Remove Claims
{num_true_fact} ground-truth facts:
@@
Appendix C Embedding Classifier Implementation Details
@@
The unused tokens ’ hidden state is used as input features for a 2-layer MLP, which is optimized with a binary cross-entropy (BCE) loss. A dropout of 0.1 is applied after the first layer. We use Adam optimizer with 1e-4 learning rate. We train each classifier for 10 epochs with batch size 8. The best validation checkpoint is saved and its performance is reported on ELOQ-Gold.
Appendix D Example News